# Cosmological Singularity and Power-Law Solutions in Modified Gravity

### v2

@@ I Introduction
To address the singularity issue, we initially explore modified theories of gravity, evaluating their viability for power-law solutions. Specifically, we examine ekpyrosis [1, 2] and accelerated power-law solutions, representing alternative cosmological scenarios to inflation and a cosmological constant, respectively. These solutions are tested for both the strong energy and null energy conditions via the Raychaudhuri equation.
@@
To study the existence of singularities in general and in particular at the big-bang, we use the Raychaudhuri equation, which as we know, governs the evolution of ‘expansion of geodesics’, and is the primary tool behind the celebrated Hawking-Penrose singularity theorems [3, 4, 5, 6, 7]. The equation predicts the existence of conjugate points, both in the finite past and the finite future. A congruence of geodesics meet at these points, making the tangent vectors ambiguous. This in turn implies the incompleteness of geodesics via a set of straightforward arguments, which is generally accepted as the necessary and sufficient condition for a spacetime to be singular [8]. The only assumption that one needs to make is that the energy-momentum under question satisfies reasonable energy conditions. These can of course be broken under certain circumstances, for example when one has a positive cosmological constant (which has negative pressure) or when certain modified theories of gravity are considered instead of GR [nojiri2011unified, nojiri2017modified].
@@ II Cosmology and Power-law Solutions
Our objective is to examine whether various modified gravity models result in cosmological power-law solutions, when avoiding the singularity. These solutions serve as cosmological attractors [9]. For example, an exponential potential yields a scale factor with a power-law behavior. Our investigation covers both late-time accelerating solutions and ekpyrotic solutions for the early universe. For power-law solutions, the scale factor is as follows:
@@
Inflation proposes that the universe underwent significant acceleration during its early stages, rapidly expanding from subatomic to macroscopic scales within a fraction of a second [10, 11]. However, the adoption of accelerating power-law solutions as a model for inflation is disregarded due to their large tensor-to-scalar ratio, as initially identified in [12], which conflicts with the Planck data. Consequently, the accelerating scenario becomes a feasible option solely for late-time cosmology.
Ekpyrotic or cyclic cosmology emerges as a popular alternative to the theory of cosmic inflation. It posits that our universe undergoes an infinite number of cycles, each beginning with a big bang phase. This is followed by a slow accelerating expansion phase, then a slow contraction phase, ultimately culminating into a big crunch. The power-law behavior becomes evident during the slow contraction phase of the ekpyrotic universe’s evolution [1, 2]. Therefore, searching for ekpyrotic solutions is mathematically very similar to the search for accelerating power-law solutions.
@@ III Raychaudhuri equation and the focusing Theorem
Geodesic congruences with time-like and null tangent vector fields are characterized by their expansion $\theta=\nabla_{\mu}u^{\mu}$, $\hat{\theta}=\nabla_{\nu}n^{\nu}$, respectively[14], which satisfy the Raychaudhuri equations
@@
When $d\theta/d\tau\leq 0$ the time-like geodesics converge to a focal point [13, 3, 15, 4, 5, 16]. Since $\theta^{2}$ is positive, for $D=4$ the convergence condition simplifies to proving only
@@
Therefore, ordinary matter such as radiation ($w=\frac{1}{3}$) and dust ($w=0$) obey (9), however, a positive cosmological constant $\Lambda$ ($w=-1$) is a classic example that violates the strong energy condition.
@@
Next, we will consider the application of three modified gravity models: Stelle gravity, $f(R)$ gravity, and the Brans-Dicke theory cosmological spacetimes. We will investigate whether we can obtain interesting power-law solutions that satisfy either the timelike or null energy condition while potentially avoiding the big-bang singularity.
@@ IV Stelle Gravity
We will start with the Stelle gravity theory that introduces higher curvature terms in an attempt to smoothen the curvature singularities at the origin. In addition to the massless graviton, it gives rise to massive spin $2$ excitations, with the action given by [17, 18, 19]
@@ IV.1 Time-like Raychaudhuri Equation:
It can be shown (see [20]) that the (00) component of the tensor $H_{\mu\nu}$ is given by
@@ IV.2 The Null Raychaudhuri Equation
The null Raychaudhuri equation on the other hand takes the form {dmath} d^θdλ=-κ^2ρ_0(w+1)t^-3α(w+1)-2α -α22a02t^-2(α+1)-12λt^-2(2+α)-3λt^-4(1+α)+t^-2(1+2α). Since $\lambda>0$, there is only one positive term on the right-hand side of (IV.2). However, at very early times, this positive term will not dominate the fourth term, which has the highest negative power of $t$, thereby leaving us with no chance of violating the convergence theorem. In principle, Stelle gravity does not violate singularity theorems, except in cases where $\lambda<0$, which introduces a tachyon into the theory, and that is clearly undesirable [21].
@@
The second model we consider is konwn as $f(R)$ gravity, which generalizes the Einstein-Hilbert action of General Relativity by replacing the Ricci scalar $R$ with a function of the Ricci scalar [22], denoted as $f(R)$. The action of $f(R)$ gravity is given by
@@ V.1 The $R^{n}$ Model
In this paper we will consider an $f(R)$ theory consisting of the usual scalar curvature plus a perturbative correction of higher power in the curvature as $f(R)=R[1+(\ell_{Pl}^{2}R)^{n-1)}]$,  [^1] Note the presence of the $l_{Pl}^{2}$ term to ensure the correct dimensions.The Planck length $\ell_{Pl}$ is controlling the strength of the corrections, and $n$ is a positive integer. Since we are interested in the large curvature regime $R\gg 1/\ell^{2}_{Pl}\implies R\ell^{2}_{Pl}\gg 1$, we will consider the case where the second term in $f(R)$ dominates so that $f(R)\simeq R^{n}$. In this paper, this is the $f(R)$ gravity model [23, 24, 25, 26, 27] that we will investigate for singularity avoidance. In the following sections, we will examine the time-like and the null Raychaudhuri equations.
@@ V.2 Time-like Raychaudhuri Equation
The time-like Raychaudhuri equation in $f(R)=R^{n}$ gravity [29] is given by
@@
$-\frac{\kappa^{2}T_{\mu\nu}u^{\mu}u^{\nu}}{2n\ell_{Pl}^{2n-2}R^{n-1}}=-\frac{\kappa^{2}}{2}T_{\mu\nu}u^{\mu}u^{\nu},$ (28)
@@
$-\frac{\theta^{2}}{3}+\frac{3}{n}\left(\frac{\ddot{a}}{a}+3\frac{\dot{a}^{2}}{a^{2}}\right)=3\frac{\ddot{a}}{a}$ (29)
@@
Now, to violate the convergence condition the sum of the coefficients of $t$ in (V.2) must be positive:
@@
$03\alpha^{2}\left(\frac{2}{n}-1\right)-\frac{3\alpha}{n}>$ 0 . (31)
@@ V.3 The Null Raychaudhuri Equation
The null Raychaudhuri equation in $f(R)=R^{n}$ gravity takes the following form [29],
@@
$-\frac{\kappa^{2}T_{\mu\nu}n^{\mu}n^{\nu}}{2n\ell_{Pl}^{2\left(n-1\right)}R^{n-1}}\simeq-\frac{\kappa^{2}}{2}T_{\mu\nu}n^{\mu}n^{\nu},$ (33)
@@
$+\frac{1}{R^{n-1}}\left(n^{t}\right)^{2}\partial^{2}_{t}R^{n-1}=0$ (34)
will not exist. The expression (32) above can thus be expressed as {dmath} d^θdλ =-12α^2t^-2α-2+(2-2n)(1-2n)nt^-2α-2. Just as in the time-like case, the time dependence on the right-hand side of the above equation is identical for all terms. Ignoring the Planck-suppressed term, the condition for violating the focusing theorem simplifies to:
@@ VI Brans-Dicke theory
The last model of interest is the Brans-Dicke theory of gravitation, which incorporates an additional long-range scalar field alongside the metric and is considered a viable alternative to General Relativity. The action for the Brans-Dicke theory in the Jordan frame is given by [30, 31]
@@
where $\mathcal{L}_{M}$ is the Lagrangian density of matter, $\phi$ is the scalar field of the theory, $g$ is the determinant of the metric tensor $g_{\mu\nu}$ and $\omega$ is the Brans-Dicke coupling constant. By varying (36) with respect to the metric tensor, we get the field equations of motion {dmath} R_μν -12R g_μν = 8πϕ T_μν+1ϕ(∇_μ∇_νϕ-g_μνϕ)+ωϕ2(∇_μϕ∇_νϕ-12g_μν∇_cϕ∇^cϕ). In addition, the variation of the action with respect to the scalar field results in the equation of motion for $\phi$
@@
when we compare the first term on the right hand side of (VI) to the source term in the usual Einstein field equations of GR. The relation between $\phi$ and G in (38) indicates that G varies with time. The coupling constant $\omega$ is a tunable parameter that can be adjusted to fit the data at hand.
@@ VI.1 Dynamics from the Friedmann Equations
The (00) component of the Einstein-Brans-Dicke equation (VI) for perfect fluid takes the form:
@@ VI.2 Time-like Raychaudhuri Equation
The curvature term of the Brans-Dicke theory (when $k=0$) is given by [32]
@@
Note that Eqs. (55) and (57) are identical. In order to violate the convergence condition, $\omega$ needs to obey either case 1, case 2, or case 3. The bounds on $\omega$ in case 1 and case 2 were discussed in [33], where the latter case was shown to model accelerated expansion in Brans-Dicke theory and the former was mentioned to be consistent with a radiation-dominated epoch of the theory. Note that when $\gamma=0$, $\phi$ becomes a constant, and the power-law solution takes the form $a(t)\propto t^{\frac{1}{\epsilon}}$, thus recovering GR.
@@
Figure 7: The left hand side (red) and right hand side (cyan) of Eq.(50) for $-1\leq\gamma\leq 1$ and a fixed value of $\omega=1$ that satisfies the inequality $\omega>0$.
In figure 4 we display the LHS of the inequality (60) for a fixed value of $\omega=-3.5/2$, which is consistent with the other condition 54) and for various values of $\alpha=1.2,2,3$ which are larger than 1. We see that there is an interval of values of $\gamma$ for which the expression in (60 )is positive, hence satisfies the inequality.
@@
Case 3: Just like in the previous case, when we use (45), we arrive at the same condition as in (60) and find that there is no interval of values for $\gamma$ in which the expression in (60) is positive; thus, the inequality is not satisfied. In Fig. 7, we plotted the LHS (red) and RHS (cyan) of (50) in the interval $-1\leq\gamma\leq 1$, with a fixed value of $\omega=1$. It is evident that the RHS is greater than the LHS, indicating the absence of accelerating solutions.
Figure 8: Displays the left hand side of the inequality Eq.(58) for various values of $-\frac{3}{2}<\omega<0$ consistent with Case 1. The blue graph corresponds to $\omega=-1$, green $\omega=-1.1$ and red $\omega=-1.2$.
Ekpyrosis in Brans-Dicke Theory: The ekpyrotic or cyclic cosmology stands out as one of the most extensively discussed alternative to the theory of cosmic inflation [10, 11]. According to this idea, the ongoing expansion of our universe is just one occurrence in an infinite series of cycles. Each cycle commences with a big bang phase, undergoes a gradual acceleration, transitions into a slow contraction phase, and concludes with a big crunch. In this study, we have also investigated these ekpyrotic solutions. Searching for ekpyrotic solutions is mathematically very similar to the search for accelerating power-law solutions. To get ekpyrosis we need to satisfy the following condition coming from Eq.(45),
@@
Figure 9: Displays the left hand side of the inequality Eq.(64) for various values of $\omega<-3/2$ (consistent with 54).
@@
As in the accelerating case, the largest possible value of $\gamma$ is 2. Now we will plot (Fig. 9) the LHS of the inequality 64 for $\omega<-3/2$ to see if there is any $\gamma$ that satisfies (64).
@@
Then, equation 64 reduces to $F(\gamma)<0$. A plot of $F(\gamma)$ (see Fig. 11) when $\omega=1,2,3$ and $-2\leq\gamma\leq 2$ reveals that $F(\gamma)$ cannot be negative. This implies that for values of $\gamma$ in the interval $-2\leq\gamma\leq 2$, equation (64) cannot hold, and thus, there is no ekpyrosis for this case.
@@
Figure 12: Displays the left hand side of the inequality Eq.(62) when $\omega=-1$ and $\alpha=0.25$ (blue), $\omega=-1.1$ and $\alpha=0.125$ (red) and $\omega=-1.2$ and $\alpha=0.0625$ (green) which are consistent with $-\frac{3}{2}<\omega<0$ in Case 1.

### v3

@@ I Introduction
To address the singularity issue, we initially explore modified theories of gravity, evaluating their viability for power-law solutions. Specifically, we examine ekpyrosis [khoury2001ekpyrotic, steinhardt2002cyclic] and accelerated power-law solutions, representing alternative cosmological scenarios to inflation and a cosmological constant, respectively. These solutions are tested for both the strong energy and null energy conditions via the Raychaudhuri equation.
@@
To study the existence of singularities in general and in particular at the big-bang, we use the Raychaudhuri equation, which as we know, governs the evolution of ‘expansion of geodesics’, and is the primary tool behind the celebrated Hawking-Penrose singularity theorems [hawking1970singularities, penrose1965gravitational, hawking2023large, hawking2010nature, ellis1977singular]. The equation predicts the existence of conjugate points, both in the finite past and the finite future. A congruence of geodesics meet at these points, making the tangent vectors ambiguous. This in turn implies the incompleteness of geodesics via a set of straightforward arguments, which is generally accepted as the necessary and sufficient condition for a spacetime to be singular [senovilla1998singularity]. The only assumption that one needs to make is that the energy-momentum under question satisfies reasonable energy conditions. These can of course be broken under certain circumstances, for example when one has a positive cosmological constant (which has negative pressure) or when certain modified theories of gravity are considered instead of GR [nojiri2011unified, nojiri2017modified].
@@ II Cosmology and Power-law Solutions
Our objective is to examine whether various modified gravity models result in cosmological power-law solutions, when avoiding the singularity. These solutions serve as cosmological attractors [ali2014power]. For example, an exponential potential yields a scale factor with a power-law behavior. Our investigation covers both late-time accelerating solutions and ekpyrotic solutions for the early universe. For power-law solutions, the scale factor is as follows:
@@
Inflation proposes that the universe underwent significant acceleration during its early stages, rapidly expanding from subatomic to macroscopic scales within a fraction of a second [guth1981inflationary, linde1982new]. However, the adoption of accelerating power-law solutions as a model for inflation is disregarded due to their large tensor-to-scalar ratio, as initially identified in [Souradeep:1992sm], which conflicts with the Planck data. Consequently, the accelerating scenario becomes a feasible option solely for late-time cosmology.
Ekpyrotic or cyclic cosmology emerges as a popular alternative to the theory of cosmic inflation. It posits that our universe undergoes an infinite number of cycles, each beginning with a big bang phase. This is followed by a slow accelerating expansion phase, then a slow contraction phase, ultimately culminating into a big crunch. The power-law behavior becomes evident during the slow contraction phase of the ekpyrotic universe’s evolution [khoury2001ekpyrotic, steinhardt2002cyclic]. Therefore, searching for ekpyrotic solutions is mathematically very similar to the search for accelerating power-law solutions.
@@ III Raychaudhuri equation and the focusing Theorem
Geodesic congruences with time-like and null tangent vector fields are characterized by their expansion $\theta=\nabla_{\mu}u^{\mu}$, $\hat{\theta}=\nabla_{\nu}n^{\nu}$, respectively[haque2017consistent], which satisfy the Raychaudhuri equations
@@
When $d\theta/d\tau\leq 0$ the time-like geodesics converge to a focal point [raychaudhuri1955relativistic, hawking1970singularities, ellis2003closed, penrose1965gravitational, hawking2023large, das2019constraints]. Since $\theta^{2}$ is positive, for $D=4$ the convergence condition simplifies to proving only
@@
Therefore, ordinary matter such as radiation ($w=\frac{1}{3}$) and dust ($w=0$) obey (9), however, a positive cosmological constant $\Lambda$ (w=1w=-1) is a classic example that violates the strong energy condition.
@@
Next, we will consider the application of three modified gravity models: Stelle gravity, $f(R)$ gravity, and the Brans-Dicke theory. We will investigate whether we can obtain interesting power-law solutions that satisfy either the timelike or null energy condition while potentially avoiding the big-bang singularity.
@@ IV Stelle Gravity
We will start with the Stelle gravity theory that introduces higher curvature terms in an attempt to smoothen the curvature singularities at the origin. In addition to the massless graviton, it gives rise to massive spin $2$ excitations, with the action given by [stelle1978classical, stelle1977renormalization, noakes1983initial]
@@ IV.1 Time-like Raychaudhuri Equation:
It can be shown (see [nenmeli2023stelle]) that the (00) component of the tensor $H_{\mu\nu}$ is given by
@@ IV.2 The Null Raychaudhuri Equation
The null Raychaudhuri equation on the other hand takes the form
$\frac{d\hat{\theta}}{d\lambda}=-\kappa^{2}\rho_{0}\left(w+1\right)t^{-3\alpha\left(w+1\right)-2\alpha}-\frac{\alpha^{2}}{2a_{0}^{2}}t^{-2(\alpha+1)}-12\lambda t^{-2(2+\alpha)}-3\lambda t^{-4(1+\alpha)}+t^{-2(1+2\alpha)}.$ (23)
Since $\lambda>0$, there is only one positive term on the right-hand side of (23). However, at very early times, this positive term will not dominate the fourth term, which has the highest negative power of $t$, thereby leaving us with no chance of violating the convergence theorem. In principle, Stelle gravity does not violate singularity theorems, except in cases where $\lambda<0$, which introduces a tachyon into the theory, and that is clearly undesirable [platania2020non].
@@
The second model we consider is konwn as $f(R)$ gravity, which generalizes the Einstein-Hilbert action of General Relativity by replacing the Ricci scalar $R$ with a function of the Ricci scalar [buchdahl1970non], denoted as $f(R)$. The action of $f(R)$ gravity is given by
@@ V.1 The $R^{n}$ Model
In this paper we will consider an $f(R)$ theory consisting of the usual scalar curvature plus a perturbative correction of higher power in the curvature as $f(R)=R[1+(\ell_{Pl}^{2}R)^{n-1)}]$,  [^1] Note the presence of the $l_{Pl}^{2}$ term to ensure the correct dimensions.The Planck length $\ell_{Pl}$ is controlling the strength of the corrections, and $n$ is a positive integer. Since we are interested in the large curvature regime $R\gg 1/\ell^{2}_{Pl}\implies R\ell^{2}_{Pl}\gg 1$, we will consider the case where the second term in $f(R)$ dominates so that $f(R)\simeq R^{n}$. In this paper, this is the $f(R)$ gravity model [starobinsky1980new, carloni2005cosmological, leach2006shear, de2016theoretical, carloni2006bounce] that we will investigate for singularity avoidance. In the following sections, we will examine the time-like and the null Raychaudhuri equations.
@@ V.2 Time-like Raychaudhuri Equation
The time-like Raychaudhuri equation in $f(R)=R^{n}$ gravity [burger2018towards] is given by
@@
$\frac{\kappa^{2}T_{\mu\nu}u^{\mu}u^{\nu}}{2n\ell_{Pl}^{2n-2}R^{n-1}}=-\frac{\kappa^{2}}{2}T_{\mu\nu}u^{\mu}u^{\nu},$ (29)
@@
$\frac{\theta^{2}}{3}+\frac{3}{n}\left(\frac{\ddot{a}}{a}+3\frac{\dot{a}^{2}}{a^{2}}\right)=3\frac{\ddot{a}}{a}$ (30)
@@
$\frac{d\theta}{d\tau}=-3\alpha^{2}t^{-2}+\frac{3}{n}\left(2\alpha^{2}-\alpha\right)t^{-2}.$ (31)
This gives rise to the necessary positive term on the right-hand side of the Raychaudhuri equation. Therefore, extending the model from Einstein gravity to $f(R)$ gravity accomplishes two things: Firstly, it makes the stress-energy term, which appears as negative, subleading, and secondly, it introduces the necessary positive term, which is not present in the Einstein case.
Now, to violate the convergence condition the sum of the coefficients of $t$ in (31) must be positive:
@@
$3\alpha^{2}\left(\frac{2}{n}-1\right)-\frac{3\alpha}{n}>0.$ (33)
@@ V.3 The Null Raychaudhuri Equation
The null Raychaudhuri equation in $f(R)=R^{n}$ gravity takes the following form [burger2018towards],
@@
$\frac{\kappa^{2}T_{\mu\nu}n^{\mu}n^{\nu}}{2n\ell_{Pl}^{2\left(n-1\right)}R^{n-1}}\simeq-\frac{\kappa^{2}}{2}T_{\mu\nu}n^{\mu}n^{\nu},$ (35)
@@
$\frac{1}{R^{n-1}}\left(n^{t}\right)^{2}\partial^{2}_{t}R^{n-1}=0$ (36)
will not exist. The expression (34) above can thus be expressed as
$\frac{d\hat{\theta}}{d\lambda}=-\frac{1}{2}\alpha^{2}t^{-2\alpha-2}+\frac{\left(2-2n\right)\left(1-2n\right)}{n}t^{-2\alpha-2}.$ (37)
Just as in the time-like case, the time dependence on the right-hand side of the above equation is identical for all terms. Ignoring the Planck-suppressed term, the condition for violating the focusing theorem simplifies to:
@@ VI Brans-Dicke theory
The last model of interest is the Brans-Dicke theory of gravitation, which incorporates an additional long-range scalar field alongside the metric and is considered a viable alternative to General Relativity. The action for the Brans-Dicke theory in the Jordan frame is given by [brans1961mach, weinberg2004gravitation]
@@
where $\mathcal{L}_{M}$ is the Lagrangian density of matter, $\phi$ is the scalar field of the theory, $g$ is the determinant of the metric tensor $g_{\mu\nu}$ and $\omega$ is the Brans-Dicke coupling constant. By varying (39) with respect to the metric tensor, we get the field equations of motion
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=\frac{8\pi}{\phi}T_{\mu\nu}+\frac{1}{\phi}\left(\nabla_{\mu}\nabla_{\nu}\phi-g_{\mu\nu}\Box\phi\right)+\frac{\omega}{\phi^{2}}\left(\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}\nabla_{c}\phi\nabla^{c}\phi\right).$ (40)
In addition, the variation of the action with respect to the scalar field results in the equation of motion for $\phi$
@@
when we compare the first term on the right hand side of (40) to the source term in the usual Einstein field equations of GR. The relation between $\phi$ and G in (42) indicates that G varies with time. The coupling constant $\omega$ is a tunable parameter that can be adjusted to fit the data at hand.
@@ VI.1 Dynamics from the Friedmann Equations
The (00) component of the Einstein-Brans-Dicke equation (40) for perfect fluid takes the form:
@@ VI.2 Time-like Raychaudhuri Equation
The curvature term of the Brans-Dicke theory (when $k=0$) is given by [choudhury2021raychaudhuri]
@@
Note that Eqs. (59) and (61) are identical. In order to violate the convergence condition, $\omega$ needs to obey either case 1, case 2, or case 3. The bounds on $\omega$ in case 1 and case 2 were discussed in [banerjee2001cosmic], where the latter case was shown to model accelerated expansion in Brans-Dicke theory and the former was mentioned to be consistent with a radiation-dominated epoch of the theory. Note that when $\gamma=0$, $\phi$ becomes a constant, and the power-law solution takes the form $a(t)\propto t^{\frac{1}{\epsilon}}$, thus recovering GR.
@@
Figure 7: The left hand side (red) and right hand side (cyan) of Eq.(54) for $1\leq\gamma\leq 1$ and a fixed value of $\omega=1$ that satisfies the inequality $\omega>0$.
In figure 4 we display the LHS of the inequality (64) for a fixed value of $\omega=-3.5/2$, which is consistent with the other condition 58) and for various values of $\alpha=1.2,2,3$ which are larger than 1. We see that there is an interval of values of $\gamma$ for which the expression in (64 )is positive, hence satisfies the inequality.
@@
Case 3: Just like in the previous case, when we use (49), we arrive at the same condition as in (64) and find that there is no interval of values for $\gamma$ in which the expression in (64) is positive; thus, the inequality is not satisfied. In Fig. 7, we plotted the LHS (red) and RHS (cyan) of (54) in the interval $1\leq\gamma\leq 1$, with a fixed value of $\omega=1$. It is evident that the RHS is greater than the LHS, indicating the absence of accelerating solutions.
Figure 8: Displays the left hand side of the inequality Eq.(62) for various values of $\frac{3}{2}<\omega<0$ consistent with Case 1. The blue graph corresponds to $\omega=-1$, green $\omega=-1.1$ and red $\omega=-1.2$.
Ekpyrosis in Brans-Dicke Theory: The ekpyrotic or cyclic cosmology stands out as one of the most extensively discussed alternative to the theory of cosmic inflation [guth1981inflationary, linde1982new]. According to this idea, the ongoing expansion of our universe is just one occurrence in an infinite series of cycles. Each cycle commences with a big bang phase, undergoes a gradual acceleration, transitions into a slow contraction phase, and concludes with a big crunch. In this study, we have also investigated these ekpyrotic solutions. Searching for ekpyrotic solutions is mathematically very similar to the search for accelerating power-law solutions. To get ekpyrosis we need to satisfy the following condition coming from Eq.(49),
@@
Figure 9: Displays the left hand side of the inequality Eq.(68) for various values of $\omega<-3/2$ (consistent with 58).
@@
As in the accelerating case, the largest possible value of $\gamma$ is 2. Now we will plot (Fig. 9) the LHS of the inequality 68 for $\omega<-3/2$ to see if there is any $\gamma$ that satisfies (68).
@@
Then, equation 68 reduces to $F(\gamma)<0$. A plot of $F(\gamma)$ (see Fig. 11) when $\omega=1,2,3$ and $2\leq\gamma\leq 2$ reveals that $F(\gamma)$ cannot be negative. This implies that for values of $\gamma$ in the interval $2\leq\gamma\leq 2$, equation (68) cannot hold, and thus, there is no ekpyrosis for this case.
@@
Figure 12: Displays the left hand side of the inequality Eq.(66) when $\omega=-1$ and $\alpha=0.25$ (blue), $\omega=-1.1$ and $\alpha=0.125$ (red) and $\omega=-1.2$ and $\alpha=0.0625$ (green) which are consistent with $\frac{3}{2}<\omega<0$ in Case 1.